### v2

# Phase-Space Topology in a Single-Atom Synthetic Dimension

Abstract
We investigate topological features in the synthetic Fock-state lattice of a single-atom system described by the quantum Rabi model. By diagonalizing the Hamiltonian, we identify a zero-energy defect state localized at a domain wall of the synthetic lattice, whose spin polarization is topologically protected. To address the challenge of applying band topology to the Fock-state lattice, we introduce a topological invariant based on phase-space geometry—the phase-space winding number. We show that the Zak phase, representing the geometric phase difference between two sublattices, can also be computed using a phase-space parameter and corresponds directly to the phase-space winding number. This quantized geometric phase reflects the spin polarization of the defect state, demonstrating a bulk–boundary correspondence. The resulting phase-space topology reveals the emergence of single-atom dressed states with contrasting properties—topologically protected fermionic states and driving-tunable bosonic states. Our results establish phase-space topology as a novel framework for exploring topological physics in single-atom synthetic dimensions, uncovering quantum-unique topological protection distinct from classical analogs.
I Introduction
A single two-level atom interacting with a harmonic oscillator, as described by the quantum Rabi model (QRM) [1] or the Jaynes–Cummings model [2], has been a prominent subject of theoretical and experimental studies [3, 4]. Along with the quasi-exact solvability of the QRM [5, 6], various theoretical [7] and experimental [8, 9] studies have explored the quantum phase transition in the QRM. Subsequent studies have investigated variants of the QRM incorporating anisotropy [10], dissipation [11], nonlinearity [12], and finite-frequency effects [13], each revealing distinct physical phenomena. In parallel, recent experimental advances have enabled the realization of ultra-strong [14, 15, 16] and deep-strong coupling regimes [17, 18, 19, 20, 21] in the QRM [4], allowing the simulation of exotic phenomena emerging from strong atom–oscillator coupling.
Since the discovery of the quantum Hall effect [22], topological phases of matter have been extensively studied across a variety of physical platforms, including optical lattices [23, 24], photonics [25, 26], trapped ions [27, 28], superconducting circuits [29, 30], and neutral atoms [31, 32]. One of the emerging topics in this field is to implement fertile physics of high-dimensional topological phenomena within physically accessible platforms, for example, using the concept of synthetic dimensions [33, 34, 35, 36, 37]. Although most of synthetic-dimensional physics have been studied in classical platforms, the use of Fock basis in quantum systems, which enables an infinite number of lattice sites even with single-mode bosonic excitations, has recently been proposed [38, 39]. For example, a one-dimensional Fock-state lattice, induced by the interaction term in the QRM [39], can be viewed as an infinite Su–Schrieffer–Heeger (SSH) chain [40]. However, due to the absence of translational symmetry, and thus, the consequent difficulties in defining quasi-momentum, Brillouin zone, and the closed form of topological invariants, the analytical estimation of topological phenomena in Fock-state lattices still remains a challenge.
In this paper, we propose a phase-space analysis of topological Fock-state lattices, which enables the characterization of well-defined topological invariants for a single-atom QRM. We derive analytic solutions and identify a zero-energy defect state localized at a domain wall, whose spin polarization indicates distinct topological phases. Furthermore, we define a phase-space winding number as a topological invariant capturing the topological protection in fermionic states in the QRM. The phase-space winding number corresponds to the quantized geometric phase difference between the two sublattices, reflecting the bulk–boundary correspondence in the system. Our results provide a novel framework for exploring topological features in a single-atom system, with potential realizations in trapped ions [19, 41] and superconducting circuits [42, 43, 44, 45].
II Defect states and energy spectrum
We consider a generalized quantum Rabi Hamiltonian [46], neglecting on-site energy terms:
[…]
where h.c. denotes the Hermitian conjugate. $\hat{\sigma}_{x}$ is the Pauli $x$-operator, $\hat{a}^{\dagger}$ ($\hat{a}$) is a bosonic creation (annihilation) operator, and operators $\hat{\sigma}_{\pm}$ denote the atomic raising and lowering operators acting on a two-level system. We treat both $v_{r}=|v_{r}|e^{i\phi_{r}}$ and $v_{cr}=|v_{cr}|e^{i\phi_{cr}}$ as complex parameters, while $w$ is assumed to be real and positive. The amplitudes $|v_{r}|$ and $|v_{cr}|$ represent the coupling strengths between the atom and the red- and blue-sideband transitions, respectively, with the phases $\phi_{r}$ and $\phi_{cr}$ denoting their relative phase with respect to the carrier drive $w$ [19, 42, 43]. The Hamiltonian possesses chiral symmetry with respect to the Pauli $z$-operator, $\hat{\sigma}_{z}$, and can be implemented in trapped-ion [19, 41] and superconducting-circuit [42, 43, 44, 45] platforms.
Figure 1: Illustration of a domain wall formed when $w\neq 0$. The upper and lower lattices correspond to the cases $v_{r}=0$ and $v_{cr}=0$, respectively. The non-zero hopping term is denoted as $v$. Each site in the synthetic lattice is labeled by a Fock state $|n\rangle$, with blue and orange circles representing the spin states $|e\rangle$ and $|g\rangle$, respectively. Solid lines indicate the stronger of the intra-cell ($w$) and inter-cell ($v$) couplings, while dashed lines represent the other. Red dotted boxes outline the dimerized unit cells in each region. The domain wall is located at the site $|n\rangle$ when $|v|\sqrt{n}<w<|v|\sqrt{n+1}$. The change in dimerization at site $|n\rangle$ creates a defect, marked by a red circle.
The Hamiltonian in Eq. (1) effectively describes an SSH-like, semi-infinite chain in a synthetic Fock-state lattice when either $v_{r}=0$ or $v_{cr}=0$ (Fig. 1). The Fock states $|n\rangle$ serve as lattice sites, with the hopping amplitude between neighboring sites scaling as $\sqrt{n}$, lacking the discrete translational symmetry. Since $w$ remains constant across all lattice sites, the increasing hopping amplitude induces a domain wall at the site $|n\rangle$ where the relative strength between $w$ and $|v|\sqrt{n}$ is reversed, where $v$ denotes the non-zero hopping parameter. In this semi-infinite SSH configuration with broken translational symmetry, the domain wall results in a defect at $|n\rangle$, as illustrated in Fig. 1.
Figure 2: Energy spectrum $E^{\pm}_{n}$ of the Hamiltonian in Eq. (1) as a function of $|v_{r}|-|v_{cr}|$, with $w=5$. The spectrum is obtained via numerical diagonalization with a Fock-space cutoff $N_{\max}=3000$, for $\phi=0$ and $\phi=\pi/2$. The singular region $-0.01\leq|v_{r}|-|v_{cr}|\leq 0.01$, which is excluded from the calculation, is masked out. Each line corresponds to an energy level $E_{n}$, and the color indicates its sign. Cross markers represent theoretical predictions from Eq. (2) for $1\leq n\leq 10$, showing good agreement with the numerical results, except near the singular region. When $\phi=\pi/2$, the spectrum displays a distinct feature near the point $|v_{r}|=|v_{cr}|$, which will be discussed later.
To unveil the dynamics in this synthetic Fock-state lattice, we derive the analytical solutions of its states and spectrum for the entire regime between two extremes illustrated in Fig. 1. For a given $w$, the dynamics over the lattice, including the emergence of defect states, is governed by the one with larger amplitude between $v_{r}$ and $v_{cr}$, while the smaller one acts as a long-range hopping correction to the system [47]. By diagonalizing the Hamiltonian in Eq. (1), we obtain the analytical solution for the complete set of energy eigenstates and eigenvalues for the general case $|v_{r}|\neq|v_{cr}|$ [48]:
$|\psi_{n+1}^{\pm}\rangle$ $=\hat{S}(\xi^{\prime})\hat{D}(\alpha)\hat{R}(\theta)\left(\frac{|n{+}1\rangle|A\rangle\pm e^{i\phi_{n}}|n\rangle|B\rangle}{\sqrt{2}}\right),$
$|\psi_{0}\rangle$ $=\hat{S}(\xi)\hat{D}(\alpha)\hat{R}(\theta)|0\rangle|A\rangle,$
$E_{n}^{\pm}$ =±$=\pm\sqrt{\left|\left(\left|v_{r}\right|-\left|v_{cr}\right|\right)n\right|},$ (2)
for $n\in\{0,1,2,\dots\}$. Here, $|A\rangle=|g\rangle$, $|B\rangle=|e\rangle$ for $|v_{r}|>|v_{cr}|$, and $|A\rangle=|e\rangle$, $|B\rangle=|g\rangle$ for $|v_{r}|<|v_{cr}|$. The operators $\hat{D}(\alpha)$, $\hat{S}(\xi)$, and $\hat{R}(\theta)$ are the displacement, squeezing, and phase-shifting operators, respectively [49, 50]. The phase factors $\phi_{n}$ indicate arbitrary relative phases between the two sublattices. The relevant parameters are defined as:
$\alpha=\frac{w(|v_{cr}|e^{-i\phi}-|v_{r}|e^{i\phi})}{|v_{r}|^{2}-|v_{cr}|^{2}},\quad\tanh(2\xi^{\prime})=\frac{2|v_{r}v_{cr}|}{|v_{r}|^{2}+|v_{cr}|^{2}},$
$\tanh\xi=\frac{\min{\{|v_{r}|,|v_{cr}|\}}}{\max{\{|v_{r}|,|v_{cr}|\}}},\ \ \theta=\frac{\phi_{cr}-\phi_{r}}{2},\ \ \phi=\frac{\phi_{r}+\phi_{cr}}{2}$ (3)
where $\xi,\xi^{\prime}\in\mathbb{R}$. Figure 2 shows the energy spectra obtained via numerical diagonalization of the Hamiltonian for $\phi=0$ and $\phi=\pi/2$. The analytical expression in Eq. (2), indicated by cross markers, shows excellent agreement with the numerical results. Notably, the spectra exhibits qualitatively different behaviors near the critical point $|v_{r}|=|v_{cr}|$, depending on the values of $\phi$, which will be discussed later.
As shown in Eq. (2), the weaker of the two coupling ($v_{r}$, $v_{cr}$) modifies the phase-space structure of the eigenstates by introducing additional displacement, rotation, and squeezing. Moreover, the zero-energy defect state $|\psi_{0}\rangle$ maintains a fixed spin polarization, $\langle\psi_{0}|\hat{\sigma}_{z}|\psi_{0}\rangle$.
When $|v_{r}|=|v_{cr}|=|v|$, the energy spectrum is:
$E^{\pm}(q,\phi)=\pm\sqrt{\frac{w^{2}}{2}\left(1-\cos{2\phi}\right)+2|v|^{2}q^{2}},$ (4)
where $q$ is the position eigenvalue in the transformed phase space defined by $\hat{q}=\hat{R}(-\theta)\hat{D}(-\beta)\hat{x}\hat{D}(\beta)\hat{R}(\theta)$, with $\beta=-\frac{we^{i\phi_{r}}}{\sqrt{2}|v|}$ and $\hat{x}=\frac{1}{\sqrt{2}}\left(\hat{a}+\hat{a}^{\dagger}\right)$ [48]. The energy gap closes at $\phi=2\pi m$ for $m\in\mathbb{Z}$, forming a periodic Dirac-cone structure in a $(1\!+\!1)$-dimensional space spanned by the phase-space coordinate $q$ and the parameter $\phi$.
III Phase-space topology
To investigate topological features of our infinite and aperiodic system, which are indicated by the topologically protected spin polarization of the defect state around a periodic Dirac cone, we develop the theoretical framework of phase-space topology.
The spin polarization of the zero-energy defect state is topologically protected and robust against noise in the hopping parameters $w$, $v_{r}$, and $v_{cr}$. It depends solely on the relative magnitude between $|v_{r}|$ and $|v_{cr}|$:
$\left\langle\psi_{0}\left|\hat{\sigma}_{z}\right|\psi_{0}\right\rangle=\text{sgn}(|v_{cr}|-|v_{r}|),$ (5)
as derived from Eq. (2). Thus, the two regimes $|v_{r}|>|v_{cr}|$ and $|v_{r}|<|v_{cr}|$ correspond to distinct topological phases, characterized by opposite spin polarizations of the defect state. Since the defect is localized at the interface between two distinct domains (see Fig. 1), it can be interpreted as a boundary property of the lattice.
As Eq. (2) shows, the energy gap vanishes in the limit $n\to\infty$, defining the bulk regime of the lattice. However, the non-uniform hopping amplitudes prevent a conventional momentum-space analysis of the system. Furthermore, the semi-infinite structure of the Fock-state lattice hinders the application of real-space topological invariants developed for systems lacking translational symmetry [51, 52, 48].
To resolve the difficulty in directly applying band topology to Fock-state lattices that lack translational symmetry, we focus on the phase parameter $\phi$, which provides alternative periodicity over the phase space. We note that the eigenstates are centered along a circular trajectory given by $\alpha(\phi)=C+Re^{i\phi}$. As $\phi$ varies from 0 to $2\pi$, $\alpha(\phi)$ traces a circular loop in Fig. 3, which corresponds to a helical trajectory of the eigenstates along the $\phi$-axis. Accordingly, the regimes $|v_{r}|>|v_{cr}|$ and $|v_{r}|<|v_{cr}|$ correspond to the geometric conditions $R>|C|$ and $R<|C|$, respectively. In analogy with the winding number in the SSH model [53], we define a phase-space winding number for the SSH-like chain in the Fock-state lattice as
[…]
This quantity indicates whether the trajectory $\alpha(\phi)$ encloses the origin in phase space, where $\phi$ substitutes the role of the Brillouin zone in defining the Zak phase in 1D periodic systems. The resulting winding number exhibits a discrete jump at the critical point $|v_{r}|=|v_{cr}|$, as shown in Fig. 3.
Figure 3: Phase-space trajectories for various values of $\Delta|v|=\left||v_{r}|-|v_{cr}|\right|$ with $w=2.5$. Red solid (blue dashed) curves correspond to the regime $|v_{r}|>|v_{cr}|$ ($|v_{r}|<|v_{cr}|$). As $\Delta|v|\to 0$, both the center and the radius of the trajectories diverge, while the intersection point between the two regimes converges to the location marked by a cross. The two regimes, characterized by a distinct topological property—spin polarization—can thus be distinguished by the phase-space winding number, which indicates whether the trajectory encloses the origin.
Figure 4: Average values of the phase-space winding number $W$ and spin polarization $\langle\psi_{0}|\hat{\sigma}_{z}|\psi_{0}\rangle$ under noise in $|v_{r}|$ and $\phi_{r}$. Amplitude noise in $|v_{r}|$ is uniformly sampled from $[0,\delta v]$, and phase noise in $\phi_{r}$ from $[-\pi,\pi]$. Results are obtained via Monte Carlo sampling with 1000 realizations, using $|v_{r}|=0.5$, $|v_{cr}|=1$, and $\phi_{r}=\phi_{cr}=0$. Both observables remain robust against arbitrary phase noise in $\phi_{r}$ and amplitude noise for $\delta v\leq|v_{cr}|-|v_{r}|$.
The proposed topological invariant provides an insight into the topological properties of SSH-like lattices and their defect states described in Fig. 1. We note that the winding number $W$ is defined through the geometric properties of the bosonic mode coupled to the spin state, such that fluctuations in the Hamiltonian parameters appear as deformations in the phase-space geometry.
We compute the phase-space winding number in Eq. (6) for various values of $w$ in Fig. 4, under noise sampled uniformly from $[0,\delta v]$ for $|v_{r}|$ and from $[-\pi,\pi]$ for $\phi_{r}$. Using the parameters $|v_{r}|=0.5$, $|v_{cr}|=1$, and $\phi_{r}=\phi_{cr}=0$, we observe that both the spin polarization and the phase-space winding number $W$ remain robust against amplitude and phase noise across different values of $w$, as long as the condition $|v_{r}|<|v_{cr}|$ is preserved, which corresponds to continuous deformation to quantum state geometry. However, as $\delta v$ goes beyond the threshold $|v_{cr}|-|v_{r}|$, the robustness of both quantities breaks down due to the loss of topological protection. Since the noise acts uniformly across all lattice sites, exceeding this threshold results in a statistical mixture of distinct topological phases, which hinders the topological nature of the system. Therefore, the proposed winding number characterizes the quantum-unique topological feature of our systems—while bosonic properties, such as the expectation values of $n$ or the phase space distribution, vary under continuous parameter deformation, the proposed invariant illustrates the topological protection of fermionic states, as shown in Fig. 4.
The physical meaning of the phase-space winding number can be understood by analyzing the geometric phase accumulated along the phase parameter $\phi$. As shown in Eqs. (2) and (6), the parameter $\phi$ characterizes the periodic structure of the eigenstates $|\psi^{\pm}_{n}\rangle$ in phase space, and can be interpreted as a quasi-momentum of the SSH-like chain defined on the Fock-state lattice. Accordingly, the accumulated phase difference between the two sublattices over $\phi\in[0,2\pi]$ corresponds to the Zak phase [54] of the SSH-like chain. As $\phi$ varies from 0 to $2\pi$, each Fock state $|n\rangle$ gains a phase of $n\pi/2$ due to the phase-shifting operator $\hat{R}(\theta)$, forming a helical structure along the $\phi$-axis. The dressed eigenstates $|\psi_{n}^{\pm}\rangle$ in Eq. (2) thus imply that the $|A\rangle$ sublattice exhibits a relative phase of $\pi/2$ compared to $|B\rangle$ for all $n$. As a result, a normalized bulk state $|\psi\rangle$ acquires a quantized geometric phase in the $|A\rangle$ sublattice as $\phi$ varies from 0 to $2\pi$.
As described in Eq. (6), the fermionic spin state $|A\rangle$, which is entangled with the bosonic state $|n+1\rangle$, is determined by the relative strength between $|v_{r}|$ and $|v_{cr}|$. As a result, the sign of the phase difference between the sublattices $|e\rangle$ and $|g\rangle$ exhibits a discrete jump at the transition point $|v_{r}|=|v_{cr}|$. To analyze this behavior, we fix $\phi_{cr}$ and examine the accumulated geometric phase difference between the two sublattices for a normalized bulk state $|\psi\rangle$ as $\phi$ varies over the interval $[0,2\pi]$ [55].
[…]
which denotes the difference between phase-space Zak phase that illustrates the same quantum state geometry with the winding number in Eq. (6). The discrete change in $\Delta\gamma$ reflects the spin polarization of the SSH-like chain, illustrating the bulk–boundary correspondence in the Fock-state lattice. The winding number, defined along a single loop in $\phi$-space, captures the difference in the helical structure between $|n\rangle$ and $|n+1\rangle$, with the sign information encoded in the geometric parameters $R$ and $C$.
IV Quantum phase-space analysis
To interpret the observed topological properties in terms of quantum states, we introduce the quantum phase-space illustration of defect states. The $\phi$-dependent energy spectrum, as shown in Fig. 2 and Eq. (4), reflects the critical behavior of defect states emerging in phase space. Figure 5 illustrates the phase-space trajectories for $|v_{r}|-|v_{cr}|=\pm 0.75$, along with the Wigner functions of the zero-energy defect states at $\phi=\pi/4$. The squeezing axes are perpendicular to the black dotted line connecting the centers of the two defect states in different regions. As the difference $|v_{r}|-|v_{cr}|\to 0$, the defect states become infinitely squeezed along the squeezing axis, and both the center $C$ and the radius $R$ of the trajectory diverge. At $\phi=0$, the centers of the eigenstates in different phases coincide, as shown in Fig. 5, and the energy gap in Eq. (4) closes at the critical point $\Delta|v|=0$. In contrast, for $\phi\neq 0$, the squeezing axes remain parallel between the two regimes, but a spatial separation between the phase-space centers emerges and reaches a maximum at $\phi=\pi/2$. This separation diverges as $\Delta|v|\to 0$, resulting in a finite energy gap at the critical point. This critical behavior underlies the Dirac-cone structure shown in Eq. (4).
In contrast to classical systems, where topological phenomena emerge from the band structure, the topological features in our system are rooted in its entanglement structure and phase-space characteristics. The defect states possess topologically protected spin polarization, and the spin–boson entanglement in Eq. (2) induces a Zak phase in the single-atom quantum Rabi system. The phase-space geometry defines the corresponding topological invariant, the phase-space winding number, and the interplay between displacement and squeezing of the eigenstates leads to a conical energy spectrum. Our results highlight phase-space geometry as a powerful framework for realizing and probing topological phenomena in hybrid quantum spin–oscillator systems [56, 57], where topology arises from intrinsic quantum properties.
Figure 5: Phase-space trajectories of $\alpha(\phi)$ for $w=2.5$. The red solid (blue dashed) trajectory corresponds to $|v_{r}|=1$, $|v_{cr}|=0.25$ ($|v_{r}|=0.25$, $|v_{cr}|=1$). The two trajectories intersect at $\phi=0$, and the Wigner functions of the zero-energy defect states at $\phi=\pi/4$ are shown for each regime. Gray dotted lines denote the squeezing axes, which are perpendicular to the black dot-dashed lines connecting the phase-space centers $\alpha(\phi)$ across the two regimes. While the squeezing axes remain parallel for all $\phi$, a spatial separation between the phase-space centers emerges and reaches its maximum at $\phi=\pi/2$. This separation diverges as $|v_{r}|-|v_{cr}|\to 0$ for $\phi\neq 0$, indicating the critical behavior that underlies the conical energy spectrum in Eq. (4).
V Discussion
In this work, we have shown that a single-atom system described by the quantum Rabi model exhibits topological features characterized by its phase-space geometry. To overcome the absence of translational symmetry and the unconventional band structure of the Fock-state lattice, we introduced a phase-space winding number as a topological invariant, reflecting the quantized geometric phase. We showed that the proposed invariant describes the topologically protected fermionic spin state in the defect states of the Rabi SSH-like chain and links directly to the Zak phase of the Fock-state lattice. Furthermore, we analyzed the $\phi$-dependent dynamics in phase space, which leads to a conical energy spectrum at the critical point. Our results demonstrate topological features emerge from the intrinsic quantum nature of a single-atom synthetic lattice, establishing a compact and controllable framework for exploring topological physics. We demonstrate that topological phases, typically characterized by spatial symmetries in real-space lattices, can instead be identified through geometric features in the phase space of a synthetic Fock-state lattice.
The Hamiltonian in Eq. (1) can be implemented using trapped ions or superconducting circuits, where the coupling strengths $|v_{r}|$, $|v_{cr}|$, and $w$ correspond to sideband and carrier Rabi frequencies [19, 41, 42, 43, 44, 45]. In trapped-ion systems, typical red- and blue-sideband Rabi frequencies are on the order of tens of kHz, and $v_{r}$, $v_{cr}$, and $w$ can be independently tuned via radio-frequency (RF) amplitude modulation. We note, however, that the accessible parameter regime may be constrained by the Lamb–Dicke condition [41], which limits both the maximum achievable sideband coupling strength and the Fock-state population. In superconducting circuits, sideband Rabi frequencies can reach tens of MHz [45], allowing for faster dynamics. However, short coherence times [42] may restrict the evolution time, and the effective parameter regimes remain similar, as the topological features of the system are determined by relative coupling strengths rather than absolute values. In both platforms, the phase factors $\phi_{r}$ and $\phi_{cr}$ are tunable by controlling the relative phases of RF electronics used for sideband generation, enabling access to the desired phase-space geometry. These systems offer precise control over the drive parameters, making them suitable for the experimental demonstration of single-atom topological features via adiabatic preparation and manipulation of defect states.
An important implication of our results is the separation between the fermionic and bosonic nature of the defect state. While the spin polarization of the defect is topologically protected and robust against parameter noise, the bosonic properties, such as the occupation number $\langle n\rangle$ and the phase-space distribution, remain tunable via system parameters. This separation enables the realization of tunable, topologically protected quantum states, where the protected spin degree of freedom acts as a stable quantum memory and the bosonic mode can be exploited for control or interfacing. Such controllable yet protected quantum states may be useful in hybrid quantum computing architectures that require coherence in one degree of freedom and tunability in another [56, 57]. Furthermore, by utilizing multiple bosonic modes in hybrid quantum devices [58, 59, 60], our system can be extended to higher-dimensional synthetic lattices. These extended systems enable the exploration of novel topological phenomena arising from interactions between spin-protected defect modes. Such developments open up new possibilities for engineering topological matter, where topology originates from phase-space characteristics.
Acknowledgements.
VI Acknowledgments
S. Yu was supported by Creative-Pioneering Researchers Program through Seoul National University. The rest of authors are supported by the Institute for Information & Communications Technology Planning & Evaluation (IITP) grant (No.RS-2022-II221040), and the National Research Foundation of Korea (NRF) grant (No. RS-2020-NR049232, No. RS-2024-00442855, No. RS-2024-00413957), all of which are funded by the Korean government (MSIT).

### v4

@@
 Abstract
-We investigate topological features in the synthetic Fock-state lattice of a single-atom system described by the quantum Rabi model.
-By diagonalizing the Hamiltonian, we identify a zero-energy defect state localized at a domain wall of the synthetic lattice, whose spin polarization is topologically protected.
-To address the challenge of applying band topology to the Fock-state lattice, we introduce a topological invariant based on phase-space geometry—the phase-space winding number.
-We show that the Zak phase, representing the geometric phase difference between two sublattices, can also be computed using a phase-space parameter and corresponds directly to the phase-space winding number.
+We investigate topological features in the synthetic Fock-state lattice (FSL) of a single-atom system described by the quantum Rabi model.
+By diagonalizing the Hamiltonian, we identify a zero-energy defect state localized at a domain wall of the FSL, whose spin polarization is topologically protected.
+To address the challenge of applying band topology to the FSL, we introduce a physically motivated and directly measurable topological invariant based on phase-space geometry—the phase-space winding number.
+We show that the Zak phase, computed using a phase-space parameter, is related to the invariant.
 This quantized geometric phase reflects the spin polarization of the defect state, demonstrating a bulk–boundary correspondence.
-The resulting phase-space topology reveals the emergence of single-atom dressed states with contrasting properties—topologically protected fermionic states and driving-tunable bosonic states.
+The resulting phase-space topology reveals the emergence of single-atom dressed states with contrasting properties—topologically protected spin states and driving-tunable bosonic states.
 Our results establish phase-space topology as a novel framework for exploring topological physics in single-atom synthetic dimensions, uncovering quantum-unique topological protection distinct from classical analogs.
-I Introduction
-A single two-level atom interacting with a harmonic oscillator, as described by the quantum Rabi model (QRM) [1] or the Jaynes–Cummings model [2], has been a prominent subject of theoretical and experimental studies [3, 4].
+Introduction—A single two-level atom interacting with a harmonic oscillator, as described by the quantum Rabi model (QRM) [1] or the Jaynes–Cummings (JC) model [2], has been a prominent subject of theoretical and experimental studies [3, 4].
 Along with the quasi-exact solvability of the QRM [5, 6], various theoretical [7] and experimental [8, 9] studies have explored the quantum phase transition in the QRM.
@@
 Since the discovery of the quantum Hall effect [22], topological phases of matter have been extensively studied across a variety of physical platforms, including optical lattices [23, 24], photonics [25, 26], trapped ions [27, 28], superconducting circuits [29, 30], and neutral atoms [31, 32].
-One of the emerging topics in this field is to implement fertile physics of high-dimensional topological phenomena within physically accessible platforms, for example, using the concept of synthetic dimensions [33, 34, 35, 36, 37].
-Although most of synthetic-dimensional physics have been studied in classical platforms, the use of Fock basis in quantum systems, which enables an infinite number of lattice sites even with single-mode bosonic excitations, has recently been proposed [38, 39].
-For example, a one-dimensional Fock-state lattice, induced by the interaction term in the QRM [39], can be viewed as an infinite Su–Schrieffer–Heeger (SSH) chain [40].
-However, due to the absence of translational symmetry, and thus, the consequent difficulties in defining quasi-momentum, Brillouin zone, and the closed form of topological invariants, the analytical estimation of topological phenomena in Fock-state lattices still remains a challenge.
-In this paper, we propose a phase-space analysis of topological Fock-state lattices, which enables the characterization of well-defined topological invariants for a single-atom QRM.
-We derive analytic solutions and identify a zero-energy defect state localized at a domain wall, whose spin polarization indicates distinct topological phases.
-Furthermore, we define a phase-space winding number as a topological invariant capturing the topological protection in fermionic states in the QRM.
-The phase-space winding number corresponds to the quantized geometric phase difference between the two sublattices, reflecting the bulk–boundary correspondence in the system.
+An emerging topic is to realize higher-dimensional topological phenomena within physically accessible platforms, for example, using the concept of synthetic dimensions [33, 34, 35, 36, 37].
+While most studies of synthetic-dimensional physics have focused on classical platforms, using the Fock basis in quantum systems to form Fock-state lattices (FSLs) has recently been investigated [38, 39, 40, 41, 42].
+In the absence of spin coupling, the QRM can be viewed as a semi-infinite one-dimensional (1D) FSL [39], and a relation to an infinite Su–Schrieffer–Heeger (SSH) chain has been suggested [43, 39].
+While finite 1D FSLs, such as two-mode JC [38] and central-spin models [39], support SSH-like edge modes, the topological properties of the semi-infinite 1D FSL described by the QRM remain unexplored.
+Here, we demonstrate topological phenomena in a generalized QRM with spin coupling.
+We derive analytic solutions and identify a zero-energy defect state localized at a domain wall, whose spin polarization is topologically protected.
+To address the absence of translational symmetry in FSLs, we propose a phase-space approach that enables analytic estimation of topological features.
+We define a physically motivated topological invariant—the phase-space winding number—that captures the topological protection of spin polarization.
+We demonstrate that the spin polarization of the defect state is robust to parameter noise, enabling high-fidelity nonclassical bosonic-state generation and seemingly contradictory active quantum functionalities—noise-robust but target-selective spin control with dissipation-induced fidelity enhancement.
+We further show that the invariant corresponds to the quantized geometric-phase difference between the two sublattices, reflecting bulk–boundary correspondence.
 Our results provide a novel framework for exploring topological features in a single-atom system, with potential realizations in trapped ions [19, 44] and superconducting circuits [45, 46, 47, 48].
-II Defect states and energy spectrum
-We consider a generalized quantum Rabi Hamiltonian [46], neglecting on-site energy terms:
-[…]
+Defect states and energy spectrum—We consider a system described by the generalized quantum Rabi Hamiltonian [49], neglecting on-site energy terms; see Fig. 1(a).
+$\hat{H}=w\hat{\sigma}_{x}+\left(v_{r}\hat{a}^{\dagger}\hat{\sigma}_{-}+v_{cr}\hat{a}\hat{\sigma}_{-}+\text{h.c.}\right),$ (1)
 where h.c.
@@
 The Hamiltonian possesses chiral symmetry with respect to the Pauli $z$-operator, $\hat{\sigma}_{z}$, and can be implemented in trapped-ion [19, 44] and superconducting-circuit [45, 46, 47, 48] platforms.
-Figure 1: Illustration of a domain wall formed when $w\neq 0$.
-The upper and lower lattices correspond to the cases $v_{r}=0$ and $v_{cr}=0$, respectively.
-The non-zero hopping term is denoted as $v$.
-Each site in the synthetic lattice is labeled by a Fock state $|n\rangle$, with blue and orange circles representing the spin states $|e\rangle$ and $|g\rangle$, respectively.
-Solid lines indicate the stronger of the intra-cell ($w$) and inter-cell ($v$) couplings, while dashed lines represent the other.
-Red dotted boxes outline the dimerized unit cells in each region.
-The domain wall is located at the site $|n\rangle$ when $|v|\sqrt{n}<w<|v|\sqrt{n+1}$.
-The change in dimerization at site $|n\rangle$ creates a defect, marked by a red circle.
+Figure 1: Illustration of (a) the setup under consideration and (b) a domain wall formed for $w\neq 0$.
+(b) The upper and lower lattices correspond to $v_{r}=0$ and $v_{cr}=0$, respectively.
+Each site is labeled by a Fock state $|n\rangle$, with black and gray circles representing the spin states $|e\rangle$ and $|g\rangle$.
+Solid and dashed lines indicate the stronger and weaker of the intra-cell ($w$) and inter-cell ($v_{r}$ or $v_{cr}$) couplings, respectively.
+Red dotted boxes mark the dimerized unit cells.
+A domain wall appears at $|n\rangle$ when $|v|\sqrt{n}<w<|v|\sqrt{n+1}$, highlighted by a green circle.
 The Hamiltonian in Eq.
-(1) effectively describes an SSH-like, semi-infinite chain in a synthetic Fock-state lattice when either $v_{r}=0$ or $v_{cr}=0$ (Fig.
+(1) effectively describes an SSH-like, semi-infinite chain in a FSL when either $v_{r}=0$ or $v_{cr}=0$ (Fig.
 1).
 The Fock states $|n\rangle$ serve as lattice sites, with the hopping amplitude between neighboring sites scaling as $\sqrt{n}$, lacking the discrete translational symmetry.
-Since $w$ remains constant across all lattice sites, the increasing hopping amplitude induces a domain wall at the site $|n\rangle$ where the relative strength between $w$ and $|v|\sqrt{n}$ is reversed, where $v$ denotes the non-zero hopping parameter.
+Since $w$ remains constant across all lattice sites, the increasing hopping amplitude induces a domain wall at the site $|n\rangle$ where the relative strength between $w$ and $|v|\sqrt{n}$ is reversed.
+Here, $v$ denotes the non-zero hopping parameter.
 In this semi-infinite SSH configuration with broken translational symmetry, the domain wall results in a defect at $|n\rangle$, as illustrated in Fig.
 1.
-Figure 2: Energy spectrum $E^{\pm}_{n}$ of the Hamiltonian in Eq.
-(1) as a function of $|v_{r}|-|v_{cr}|$, with $w=5$.
-The spectrum is obtained via numerical diagonalization with a Fock-space cutoff $N_{\max}=3000$, for $\phi=0$ and $\phi=\pi/2$.
-The singular region $-0.01\leq|v_{r}|-|v_{cr}|\leq 0.01$, which is excluded from the calculation, is masked out.
-Each line corresponds to an energy level $E_{n}$, and the color indicates its sign.
-Cross markers represent theoretical predictions from Eq.
-(2) for $1\leq n\leq 10$, showing good agreement with the numerical results, except near the singular region.
-When $\phi=\pi/2$, the spectrum displays a distinct feature near the point $|v_{r}|=|v_{cr}|$, which will be discussed later.
-To unveil the dynamics in this synthetic Fock-state lattice, we derive the analytical solutions of its states and spectrum for the entire regime between two extremes illustrated in Fig.
+To unveil the dynamics in this synthetic FSL, we derive the analytical solutions of its states and spectrum for the entire regime between two extremes illustrated in Fig.
 1.
@@
 (1), we obtain the analytical solution for the complete set of energy eigenstates and eigenvalues for the general case $|v_{r}|\neq|v_{cr}|$ [51]:
-$|\psi_{n+1}^{\pm}\rangle$ $=\hat{S}(\xi^{\prime})\hat{D}(\alpha)\hat{R}(\theta)\left(\frac{|n{+}1\rangle|A\rangle\pm e^{i\phi_{n}}|n\rangle|B\rangle}{\sqrt{2}}\right),$
-$|\psi_{0}\rangle$ $=\hat{S}(\xi)\hat{D}(\alpha)\hat{R}(\theta)|0\rangle|A\rangle,$
-$E_{n}^{\pm}$ =±$=\pm\sqrt{\left|\left(\left|v_{r}\right|-\left|v_{cr}\right|\right)n\right|},$ (2)
-for $n\in\{0,1,2,\dots\}$.
-Here, $|A\rangle=|g\rangle$, $|B\rangle=|e\rangle$ for $|v_{r}|>|v_{cr}|$, and $|A\rangle=|e\rangle$, $|B\rangle=|g\rangle$ for $|v_{r}|<|v_{cr}|$.
-The operators $\hat{D}(\alpha)$, $\hat{S}(\xi)$, and $\hat{R}(\theta)$ are the displacement, squeezing, and phase-shifting operators, respectively [49, 50].
-The phase factors $\phi_{n}$ indicate arbitrary relative phases between the two sublattices.
-The relevant parameters are defined as:
-$\alpha=\frac{w(|v_{cr}|e^{-i\phi}-|v_{r}|e^{i\phi})}{|v_{r}|^{2}-|v_{cr}|^{2}},\quad\tanh(2\xi^{\prime})=\frac{2|v_{r}v_{cr}|}{|v_{r}|^{2}+|v_{cr}|^{2}},$
-$\tanh\xi=\frac{\min{\{|v_{r}|,|v_{cr}|\}}}{\max{\{|v_{r}|,|v_{cr}|\}}},\ \ \theta=\frac{\phi_{cr}-\phi_{r}}{2},\ \ \phi=\frac{\phi_{r}+\phi_{cr}}{2}$ (3)
-where $\xi,\xi^{\prime}\in\mathbb{R}$.
-Figure 2 shows the energy spectra obtained via numerical diagonalization of the Hamiltonian for $\phi=0$ and $\phi=\pi/2$.
-The analytical expression in Eq.
-(2), indicated by cross markers, shows excellent agreement with the numerical results.
-Notably, the spectra exhibits qualitatively different behaviors near the critical point $|v_{r}|=|v_{cr}|$, depending on the values of $\phi$, which will be discussed later.
+$|\psi_{0}\rangle$ $=\hat{U}|0\rangle|A\rangle,\quad|\psi_{n+1}^{\pm}\rangle=\hat{U}\left(\frac{|n{+}1\rangle|A\rangle\pm|n\rangle|B\rangle}{\sqrt{2}}\right),$
+$E_{n}^{\pm}$ $=\pm\sqrt{\left|\left(|v_{r}|^{2}-|v_{cr}|^{2}\right)n\right|},$ (2)
+for $n\in\{0,1,2,\dots\}$ and $\hat{U}=\hat{R}(\theta)\hat{D}(\alpha)\hat{S}(r)$.
+Here, $|A\rangle=|g\rangle$ for $|v_{r}|>|v_{cr}|$, $|A\rangle=|e\rangle$ for $|v_{r}|<|v_{cr}|$, and $|B\rangle=\hat{\sigma}_{x}|A\rangle$.
+The operators $\hat{D}(\alpha)$, $\hat{S}(\xi)$, and $\hat{R}(\theta)$ are the displacement, squeezing, and phase-shifting operators, respectively [52, 53], where
+$\alpha=\frac{w(v_{cr}^{*}-v_{r})}{|v_{r}|^{2}-|v_{cr}|^{2}},\quad\tanh(2r)=\frac{2|v_{r}v_{cr}|}{|v_{r}|^{2}+|v_{cr}|^{2}},$ (3)
+for $r\in\mathbb{R}$, $\theta=(\phi_{cr}-\phi_{r})/2$, $\phi=(\phi_{r}+\phi_{cr})/2$, where $(\cdot)^{*}$ denotes complex conjugation (see Appendix A for the comparison between analytic expression and numerical results).
 As shown in Eq.
-(2), the weaker of the two coupling ($v_{r}$, $v_{cr}$) modifies the phase-space structure of the eigenstates by introducing additional displacement, rotation, and squeezing.
+(2), the weaker of the two couplings ($v_{r}$, $v_{cr}$) modifies the phase-space structure of the eigenstates by introducing additional displacement, rotation, and squeezing.
 Moreover, the zero-energy defect state $|\psi_{0}\rangle$ maintains a fixed spin polarization, $\langle\psi_{0}|\hat{\sigma}_{z}|\psi_{0}\rangle$.
 When $|v_{r}|=|v_{cr}|=|v|$, the energy spectrum is:
-$E^{\pm}(q,\phi)=\pm\sqrt{\frac{w^{2}}{2}\left(1-\cos{2\phi}\right)+2|v|^{2}q^{2}},$ (4)
-where $q$ is the position eigenvalue in the transformed phase space defined by $\hat{q}=\hat{R}(-\theta)\hat{D}(-\beta)\hat{x}\hat{D}(\beta)\hat{R}(\theta)$, with $\beta=-\frac{we^{i\phi_{r}}}{\sqrt{2}|v|}$ and $\hat{x}=\frac{1}{\sqrt{2}}\left(\hat{a}+\hat{a}^{\dagger}\right)$ [48].
-The energy gap closes at $\phi=2\pi m$ for $m\in\mathbb{Z}$, forming a periodic Dirac-cone structure in a $(1\!+\!1)$-dimensional space spanned by the phase-space coordinate $q$ and the parameter $\phi$.
-III Phase-space topology
-To investigate topological features of our infinite and aperiodic system, which are indicated by the topologically protected spin polarization of the defect state around a periodic Dirac cone, we develop the theoretical framework of phase-space topology.
+$E^{\pm}(x,\phi)=\pm\sqrt{w^{2}\left(1-\cos{2\phi}\right)/2+2|v|^{2}x^{2}},$ (4)
+where $x$ is the eigenvalue of the position operator $\hat{x}=\frac{1}{\sqrt{2}}\left(\hat{a}+\hat{a}^{\dagger}\right)$ [51].
+The energy gap closes at $\phi=m\pi$ for $m\in\mathbb{Z}$, forming a periodic Dirac-cone structure in a (1+1)(1\!+\!1)-dimensional space spanned by the phase-space coordinate $x$ and the parameter $\phi$ (see Appendix).
+Phase-space topology—To investigate topological features of our semi-infinite and aperiodic system, indicated by the topologically protected spin polarization of a defect state around a periodic Dirac cone, we develop a theoretical framework of phase-space topology.
+This framework is related to prior approaches in dynamical systems theory [54], classical optics [55], and continuous systems [56, 57].
 The spin polarization of the zero-energy defect state is topologically protected and robust against noise in the hopping parameters $w$, $v_{r}$, and $v_{cr}$.
@@
 As Eq.
-(2) shows, the energy gap vanishes in the limit $n\to\infty$, defining the bulk regime of the lattice.
-However, the non-uniform hopping amplitudes prevent a conventional momentum-space analysis of the system.
-Furthermore, the semi-infinite structure of the Fock-state lattice hinders the application of real-space topological invariants developed for systems lacking translational symmetry [51, 52, 48].
-To resolve the difficulty in directly applying band topology to Fock-state lattices that lack translational symmetry, we focus on the phase parameter $\phi$, which provides alternative periodicity over the phase space.
-We note that the eigenstates are centered along a circular trajectory given by $\alpha(\phi)=C+Re^{i\phi}$.
-As $\phi$ varies from 0 to $2\pi$, $\alpha(\phi)$ traces a circular loop in Fig.
-3, which corresponds to a helical trajectory of the eigenstates along the $\phi$-axis.
+(2) shows, the difference in energy $E_{n}^{\pm}$ and $E_{n\pm 1}^{\pm}$ vanishes in the limit $n\to\infty$, defining the bulk regime of the lattice.
+In this limit, a conventional momentum-space analysis can be applied approximately, yielding a winding number in analogy with the SSH model, but it provides neither a clear physical interpretation nor an experimentally measurable quantity.
+Alternatively, one can apply real-space topological invariants [58, 59], developed for systems without translational symmetry, to FSLs.
+However, these invariants may require local measurements of Fock states $|n\rangle$ at $n$ beyond a parameter-dependent cutoff [51], which is impractical on current quantum hardware, in contrast to real-space systems.
+Figure 2: Phase-space trajectories for $\Delta|v|\equiv\bigl||v_{r}|-|v_{cr}|\bigr|$ at $w=2.5$.
+Red solid (blue dashed) curves correspond to the regime $|v_{r}|>|v_{cr}|$ ($|v_{r}|<|v_{cr}|$).
+As $\Delta|v|\to 0$, both the center and the radius of the trajectories diverge, while the intersection of the two regimes converges to the location marked by a cross.
+The phase-space winding number indicates whether the trajectory encloses the origin.
+To resolve the difficulty in directly applying band topology to FSLs that lack translational symmetry, we focus on the phase parameter $\phi$, which provides an alternative periodicity over the phase space.
+To remove the additional degree of freedom, we fix $\phi_{cr}$ and move to a frame rotated by $e^{i\phi_{cr}}$, such that the eigenstates are centered along a circular trajectory $\alpha(\phi)=C+Re^{2i\phi}$ in phase space.
+As $\phi$ varies from 0 to $\pi$, $\alpha(\phi)$ traces a circular loop in Fig.
+2, corresponding to a helical trajectory of the eigenstates along the $\theta$-axis (see Appendix B).
 Accordingly, the regimes $|v_{r}|>|v_{cr}|$ and $|v_{r}|<|v_{cr}|$ correspond to the geometric conditions $R>|C|$ and $R<|C|$, respectively.
-In analogy with the winding number in the SSH model [53], we define a phase-space winding number for the SSH-like chain in the Fock-state lattice as
-[…]
+In analogy with the winding number in the SSH model [60], we define a phase-space winding number for the SSH-like chain in the FSL as:
+$W=\frac{1}{2\pi i}\int_{0}^{\pi}d\phi\,\frac{d}{d\phi}\ln\alpha(\phi)=\frac{1-\text{sgn}(|v_{cr}|-|v_{r}|)}{2}.$ (6)
 This quantity indicates whether the trajectory $\alpha(\phi)$ encloses the origin in phase space, where $\phi$ substitutes the role of the Brillouin zone in defining the Zak phase in 1D periodic systems.
 The resulting winding number exhibits a discrete jump at the critical point $|v_{r}|=|v_{cr}|$, as shown in Fig.
-3.
-Figure 3: Phase-space trajectories for various values of $\Delta|v|=\left||v_{r}|-|v_{cr}|\right|$ with $w=2.5$.
-Red solid (blue dashed) curves correspond to the regime $|v_{r}|>|v_{cr}|$ ($|v_{r}|<|v_{cr}|$).
-As $\Delta|v|\to 0$, both the center and the radius of the trajectories diverge, while the intersection point between the two regimes converges to the location marked by a cross.
-The two regimes, characterized by a distinct topological property—spin polarization—can thus be distinguished by the phase-space winding number, which indicates whether the trajectory encloses the origin.
-Figure 4: Average values of the phase-space winding number $W$ and spin polarization $\langle\psi_{0}|\hat{\sigma}_{z}|\psi_{0}\rangle$ under noise in $|v_{r}|$ and $\phi_{r}$.
-Amplitude noise in $|v_{r}|$ is uniformly sampled from $[0,\delta v]$, and phase noise in $\phi_{r}$ from $[-\pi,\pi]$.
-Results are obtained via Monte Carlo sampling with 1000 realizations, using $|v_{r}|=0.5$, $|v_{cr}|=1$, and $\phi_{r}=\phi_{cr}=0$.
-Both observables remain robust against arbitrary phase noise in $\phi_{r}$ and amplitude noise for $\delta v\leq|v_{cr}|-|v_{r}|$.
-The proposed topological invariant provides an insight into the topological properties of SSH-like lattices and their defect states described in Fig.
-1.
-We note that the winding number $W$ is defined through the geometric properties of the bosonic mode coupled to the spin state, such that fluctuations in the Hamiltonian parameters appear as deformations in the phase-space geometry.
-We compute the phase-space winding number in Eq.
-(6) for various values of $w$ in Fig.
-4, under noise sampled uniformly from $[0,\delta v]$ for $|v_{r}|$ and from $[-\pi,\pi]$ for $\phi_{r}$.
-Using the parameters $|v_{r}|=0.5$, $|v_{cr}|=1$, and $\phi_{r}=\phi_{cr}=0$, we observe that both the spin polarization and the phase-space winding number $W$ remain robust against amplitude and phase noise across different values of $w$, as long as the condition $|v_{r}|<|v_{cr}|$ is preserved, which corresponds to continuous deformation to quantum state geometry.
-However, as $\delta v$ goes beyond the threshold $|v_{cr}|-|v_{r}|$, the robustness of both quantities breaks down due to the loss of topological protection.
-Since the noise acts uniformly across all lattice sites, exceeding this threshold results in a statistical mixture of distinct topological phases, which hinders the topological nature of the system.
-Therefore, the proposed winding number characterizes the quantum-unique topological feature of our systems—while bosonic properties, such as the expectation values of $n$ or the phase space distribution, vary under continuous parameter deformation, the proposed invariant illustrates the topological protection of fermionic states, as shown in Fig.
-4.
+2.
+See Appendix B for a detailed calculation of $W$.
+Figure 3: Noise simulation results averaged over 20 Monte Carlo trials at $w=v_{r}=1$ and $v_{cr}=0.25$.
+(a) Fidelity between the exact eigenstates and states obtained by adiabatic preparation.
+States are adiabatically prepared into $\left|\psi_{0}\right\rangle$ and $\left|\psi^{+}_{1,2,3}\right\rangle$ under noise on $v_{r}$ and $v_{cr}$ with amplitude $\delta v$.
+Solid and dashed lines show results without ($\Gamma=0$) and with ($\Gamma=0.25$) dissipation, respectively, for the pumping rate $\Gamma$.
+(b) Spin-flip operation via parameter quenching under fixed noise amplitude $\delta v=0.1$.
+The parameters $v_{r}$ and $v_{cr}$ are set to 0 to induce a spin-flip operation due to $w\hat{\sigma}_{x}$.
+After the flip, $v_{r}$ and $v_{cr}$ are restored with their magnitudes also flipped.
+The well-defined topological quantity in Eq.
+(6) implies that defect states possess robustness against a specific type of parameter noise—which corresponds to continuous deformation of the system—allowing for practical implications for quantum-state engineering.
+For instance, Fig.
+3(a) shows that the zero-energy defect state $\left|\psi_{0}\right\rangle$ at $w=v_{r}=1$ and $v_{cr}=0.25$ can be generated with high fidelity under small fluctuations in $v_{r}$ and $v_{cr}$.
+We initialize the system in the $\alpha=r=0$ state of Eq.
+(2) and adiabatically ramp $w$ and $v_{cr}$ in Eq.
+(3) from 0 to their target values, with $v_{r}$ fixed.
+During the adiabatic process, uniformly random fluctuations are added to the real and imaginary parts of $v_{r}$ and $v_{cr}$, with amplitudes bounded by $\delta v$.
+In contrast, the adiabatically prepared states $\left|\psi_{1,2,3}^{+}\right\rangle$ exhibit rapidly decaying fidelities for larger noise amplitudes due to the absence of topological features.
+Interestingly, this robustness is enhanced under the dissipation induced by a conventional Lindblad jump operator [61], $\hat{\sigma}_{-}$, which results in dissipative pumping into the $\left|g\right\rangle$ sublattice [44, 62], in sharp contrast to fidelity degradation of the other eigenstates (dashed lines for a pumping rate $\Gamma=0.25$).
+(See Appendix C and the Supplemental Material [51]).
+Along with robustness, our winding-number-based analysis, revealing the dependency on the relative magnitude of $v_{r}$ and $v_{cr}$, enables a seemingly contradictory capability: highly selective tunability.
+We note that setting $v_{r}$ and $v_{cr}$ to zero and then restoring them with flipped magnitudes induces a spin flip via the $w\,\hat{\sigma}_{x}$ interaction.
+Figure 3(b) demonstrates that only the zero-energy defect state $\left|\psi_{0}\right\rangle$ exhibits a robust, quantized change in $\langle\hat{\sigma}_{z}\rangle$, whereas other eigenstates show fluctuations due to parameter noise.
+Accordingly, the zero-energy defect state enables a noise-robust, quantized spin control.
+Moreover, it allows for direct measurement of the phase-space winding number via oscillator-state tomography; see Supplemental Material [51].
 The physical meaning of the phase-space winding number can be understood by analyzing the geometric phase accumulated along the phase parameter $\phi$.
 As shown in Eqs.
-(2) and (6), the parameter $\phi$ characterizes the periodic structure of the eigenstates $|\psi^{\pm}_{n}\rangle$ in phase space, and can be interpreted as a quasi-momentum of the SSH-like chain defined on the Fock-state lattice.
-Accordingly, the accumulated phase difference between the two sublattices over $\phi\in[0,2\pi]$ corresponds to the Zak phase [54] of the SSH-like chain.
-As $\phi$ varies from 0 to $2\pi$, each Fock state $|n\rangle$ gains a phase of $n\pi/2$ due to the phase-shifting operator $\hat{R}(\theta)$, forming a helical structure along the $\phi$-axis.
+(2) and (6), $\phi$ characterizes the periodic structure of the eigenstates $|\psi^{\pm}_{n}\rangle$ in phase space, and can be interpreted as a quasi-momentum of the SSH-like chain defined on the FSL.
+Accordingly, the accumulated phase difference between the two sublattices over $\phi\in[0,\pi]$ corresponds to the Zak phase [63] of the SSH-like chain.
+As $\phi$ varies from 0 to $\pi$, each Fock state $|n\rangle$ gains a phase proportional to $n$ due to the phase-shifting operator $\hat{R}(\theta)$, forming a helical structure along the $\phi$-axis.
 The dressed eigenstates $|\psi_{n}^{\pm}\rangle$ in Eq.
-(2) thus imply that the $|A\rangle$ sublattice exhibits a relative phase of $\pi/2$ compared to $|B\rangle$ for all $n$.
-As a result, a normalized bulk state $|\psi\rangle$ acquires a quantized geometric phase in the $|A\rangle$ sublattice as $\phi$ varies from 0 to $2\pi$.
+(2) thus exhibit a uniform relative phase difference between the $|A\rangle$ and $|B\rangle$ sublattices for all $n$.
+As a result, a normalized bulk state $|\psi\rangle$ acquires a quantized geometric phase in the $|A\rangle$ sublattice as $\phi$ varies from 0 to $\pi$ [51].
 As described in Eq.
-(6), the fermionic spin state $|A\rangle$, which is entangled with the bosonic state $|n+1\rangle$, is determined by the relative strength between $|v_{r}|$ and $|v_{cr}|$.
+(6), the spin state $|A\rangle$, which is entangled with the bosonic state $|n+1\rangle$, is determined by the relative strength between $|v_{r}|$ and $|v_{cr}|$.
 As a result, the sign of the phase difference between the sublattices $|e\rangle$ and $|g\rangle$ exhibits a discrete jump at the transition point $|v_{r}|=|v_{cr}|$.
-To analyze this behavior, we fix $\phi_{cr}$ and examine the accumulated geometric phase difference between the two sublattices for a normalized bulk state $|\psi\rangle$ as $\phi$ varies over the interval $[0,2\pi]$ [55].
-[…]
-which denotes the difference between phase-space Zak phase that illustrates the same quantum state geometry with the winding number in Eq.
+To analyze this behavior, we evaluate the accumulated geometric phase difference between the two sublattices for a normalized bulk state $|\psi\rangle$ as $\phi$ varies over the interval $0,\pi] [$64, 65, 51]:
+$\Delta\gamma=\frac{i}{\pi\cosh{2r}}\int_{0}^{\pi}d\phi\,\langle\psi|\hat{\sigma}_{z}|\partial_{\phi}\psi\rangle=\begin{cases}1&|v_{r}|>|v_{cr}|\\ -1&|v_{r}|<|v_{cr}|,\end{cases}$
+which defines the phase-space Zak phase that illustrates a quantum state geometry similar to the winding number in Eq.
 (6).
-The discrete change in $\Delta\gamma$ reflects the spin polarization of the SSH-like chain, illustrating the bulk–boundary correspondence in the Fock-state lattice.
+The discrete change in $\Delta\gamma$ at $|v_{r}|=|v_{cr}|$ reflects the spin polarization of the SSH-like chain, illustrating the bulk–boundary correspondence in the FSL.
 The winding number, defined along a single loop in $\phi$-space, captures the difference in the helical structure between $|n\rangle$ and $|n+1\rangle$, with the sign information encoded in the geometric parameters $R$ and $C$.
-IV Quantum phase-space analysis
-To interpret the observed topological properties in terms of quantum states, we introduce the quantum phase-space illustration of defect states.
-The $\phi$-dependent energy spectrum, as shown in Fig.
-2 and Eq.
-(4), reflects the critical behavior of defect states emerging in phase space.
-Figure 5 illustrates the phase-space trajectories for $|v_{r}|-|v_{cr}|=\pm 0.75$, along with the Wigner functions of the zero-energy defect states at $\phi=\pi/4$.
-The squeezing axes are perpendicular to the black dotted line connecting the centers of the two defect states in different regions.
+Figure 4: Phase-space trajectories of $\alpha(\phi)$ at $w=2.5$ in the frame rotated by $\phi_{cr}$.
+Red solid (blue dashed) trajectory corresponds to $|v_{r}|=1$, $|v_{cr}|=0.25$ ($|v_{r}|=0.25$, $|v_{cr}|=1$).
+The trajectories intersect at $\phi=0$, and the Wigner functions of the defect states at $\phi=\pi/4$ are shown for each regime.
+Gray dotted lines indicate the squeezing axes, perpendicular to the black dot–dashed lines connecting the phase-space centers $\alpha(\phi)$ between the two regimes.
+Quantum phase-space analysis—To interpret the observed topological properties in terms of quantum states, we introduce the quantum phase-space illustration of defect states.
+The $\phi$-dependent energy spectrum, as shown in Eq.
+(4) and Appendix A, reflects the critical behavior of defect states emerging in phase space.
+Figure 4 illustrates the phase-space trajectories of $\alpha(\phi)$ for $|v_{r}|-|v_{cr}|=\pm 0.75$, along with the Wigner functions $\mathcal{W}(\beta)\propto\langle\hat{D}(-\beta)\hat{\Pi}\hat{D}(\beta)\rangle$ of the zero-energy defect states at $\phi=\pi/4$, shown in the frame rotated by $\phi_{cr}$.
+Here, $\hat{\Pi}$ denotes the parity operator [66].
+The squeezing axes are perpendicular to the black dot-dashed line connecting the centers of the two defect states in different regions.
 As the difference $|v_{r}|-|v_{cr}|\to 0$, the defect states become infinitely squeezed along the squeezing axis, and both the center $C$ and the radius $R$ of the trajectory diverge.
 At $\phi=0$, the centers of the eigenstates in different phases coincide, as shown in Fig.
-5, and the energy gap in Eq.
+4, and the energy gap in Eq.
 (4) closes at the critical point $\Delta|v|=0$.
@@
 This separation diverges as $\Delta|v|\to 0$, resulting in a finite energy gap at the critical point.
-This critical behavior underlies the Dirac-cone structure shown in Eq.
-(4).
+This critical behavior underlies the Dirac-cone structure described in Eq.
+(4) and Appendix A.
 In contrast to classical systems, where topological phenomena emerge from the band structure, the topological features in our system are rooted in its entanglement structure and phase-space characteristics.
-The defect states possess topologically protected spin polarization, and the spin–boson entanglement in Eq.
+The defect state possesses topologically protected spin polarization, and the spin–boson entanglement in Eq.
 (2) induces a Zak phase in the single-atom quantum Rabi system.
@@
 Our results highlight phase-space geometry as a powerful framework for realizing and probing topological phenomena in hybrid quantum spin–oscillator systems [67, 68], where topology arises from intrinsic quantum properties.
-Figure 5: Phase-space trajectories of $\alpha(\phi)$ for $w=2.5$.
-The red solid (blue dashed) trajectory corresponds to $|v_{r}|=1$, $|v_{cr}|=0.25$ ($|v_{r}|=0.25$, $|v_{cr}|=1$).
-The two trajectories intersect at $\phi=0$, and the Wigner functions of the zero-energy defect states at $\phi=\pi/4$ are shown for each regime.
-Gray dotted lines denote the squeezing axes, which are perpendicular to the black dot-dashed lines connecting the phase-space centers $\alpha(\phi)$ across the two regimes.
-While the squeezing axes remain parallel for all $\phi$, a spatial separation between the phase-space centers emerges and reaches its maximum at $\phi=\pi/2$.
-This separation diverges as $|v_{r}|-|v_{cr}|\to 0$ for $\phi\neq 0$, indicating the critical behavior that underlies the conical energy spectrum in Eq.
-(4).
-V Discussion
-In this work, we have shown that a single-atom system described by the quantum Rabi model exhibits topological features characterized by its phase-space geometry.
-To overcome the absence of translational symmetry and the unconventional band structure of the Fock-state lattice, we introduced a phase-space winding number as a topological invariant, reflecting the quantized geometric phase.
-We showed that the proposed invariant describes the topologically protected fermionic spin state in the defect states of the Rabi SSH-like chain and links directly to the Zak phase of the Fock-state lattice.
-Furthermore, we analyzed the $\phi$-dependent dynamics in phase space, which leads to a conical energy spectrum at the critical point.
-Our results demonstrate topological features emerge from the intrinsic quantum nature of a single-atom synthetic lattice, establishing a compact and controllable framework for exploring topological physics.
-We demonstrate that topological phases, typically characterized by spatial symmetries in real-space lattices, can instead be identified through geometric features in the phase space of a synthetic Fock-state lattice.
+Discussion—In this work, we have shown that a single-atom quantum Rabi system exhibits topological features characterized by its phase-space geometry.
+To address the absence of translational symmetry and the unconventional band structure in FSLs, we introduced a phase-space winding number as a topological invariant that reflects a quantized geometric phase.
+This invariant captures the topologically protected spin polarization of defect states in a semi-infinite QRM chain and connects directly to the Zak phase of the FSL.
+We further analyzed the $\phi$-dependent dynamics in phase space, which led to a conical energy spectrum emerging at the critical point.
+Our results show that topological features can arise from the intrinsic quantum nature of a single-atom synthetic lattice, providing a compact and controllable framework for exploring topological phenomena in quantum systems.
+In particular, we showed that topological phases, typically characterized by spatial symmetries in real-space lattices, can instead be identified by geometric features in the phase space of a FSL.
+This provides a physically motivated, experimentally accessible topological quantity that bridges theoretical studies of semi-infinite FSLs with experimental realizations.
+For example, recent advances in bosonic state tomography [69, 70, 71] enable direct measurement of the phase-space winding number via tomographic reconstruction [51].
+We briefly discuss experimental implementations on trapped-ion and superconducting-circuit platforms in Appendix D and in the Supplemental Material [51].
+Another important implication of our results is the separation of spin and bosonic nature in the defect state.
+The spin polarization is topologically protected and robust to parameter noise, whereas bosonic properties, such as the occupation $\langle n\rangle$, designed selective flipping of spin polarization, and the phase-space distribution, remain tunable via system parameters.
+We demonstrated high-fidelity preparation of a nonclassical bosonic state under parameter noise and further enhanced its robustness via engineered dissipation to the $\left|g\right\rangle$ sublattice.
+These results open a pathway to the generation of precise nonclassical states even in the presence of noise, with potential applications in quantum metrology [72, 73, 74, 75, 76].
 The Hamiltonian in Eq.
-(1) can be implemented using trapped ions or superconducting circuits, where the coupling strengths $|v_{r}|$, $|v_{cr}|$, and $w$ correspond to sideband and carrier Rabi frequencies [19, 41, 42, 43, 44, 45].
-In trapped-ion systems, typical red- and blue-sideband Rabi frequencies are on the order of tens of kHz, and $v_{r}$, $v_{cr}$, and $w$ can be independently tuned via radio-frequency (RF) amplitude modulation.
-We note, however, that the accessible parameter regime may be constrained by the Lamb–Dicke condition [41], which limits both the maximum achievable sideband coupling strength and the Fock-state population.
-In superconducting circuits, sideband Rabi frequencies can reach tens of MHz [45], allowing for faster dynamics.
-However, short coherence times [42] may restrict the evolution time, and the effective parameter regimes remain similar, as the topological features of the system are determined by relative coupling strengths rather than absolute values.
-In both platforms, the phase factors $\phi_{r}$ and $\phi_{cr}$ are tunable by controlling the relative phases of RF electronics used for sideband generation, enabling access to the desired phase-space geometry.
-These systems offer precise control over the drive parameters, making them suitable for the experimental demonstration of single-atom topological features via adiabatic preparation and manipulation of defect states.
-An important implication of our results is the separation between the fermionic and bosonic nature of the defect state.
-While the spin polarization of the defect is topologically protected and robust against parameter noise, the bosonic properties, such as the occupation number $\langle n\rangle$ and the phase-space distribution, remain tunable via system parameters.
-This separation enables the realization of tunable, topologically protected quantum states, where the protected spin degree of freedom acts as a stable quantum memory and the bosonic mode can be exploited for control or interfacing.
-Such controllable yet protected quantum states may be useful in hybrid quantum computing architectures that require coherence in one degree of freedom and tunability in another [56, 57].
-Furthermore, by utilizing multiple bosonic modes in hybrid quantum devices [58, 59, 60], our system can be extended to higher-dimensional synthetic lattices.
-These extended systems enable the exploration of novel topological phenomena arising from interactions between spin-protected defect modes.
-Such developments open up new possibilities for engineering topological matter, where topology originates from phase-space characteristics.
-Acknowledgements.
-VI Acknowledgments
-S.
-Yu was supported by Creative-Pioneering Researchers Program through Seoul National University.
+(1) is widely used for gate operations in modern quantum processors [67], for example to generate spin–motion entanglement in trapped-ion systems [77].
+The same interaction enables multiqubit entanglement via common motional modes [78, 79], and gates robust to parameter noise have been developed [80, 81, 82, 83, 84].
+An alternative approach employs bosonic codes with spin-mediated interactions and readout [85, 71, 86, 87].
+We investigated distinctive phase-space dynamics and the resulting conical spectrum induced by a coherent spin drive $w\hat{\sigma}_{x}$ at the critical point $|v_{r}|=|v_{cr}|$.
+These characteristics suggest potential routes to robust gate design and bosonic entangling operations.
+We expect our results to extend to topological phenomena arising from interactions between spin-protected defect modes, connecting quantum information science and topological physics through a phase-space perspective.
+Acknowledgments—Sunkyu Yu was supported by Creative-Pioneering Researchers Program through Seoul National University.
 The rest of authors are supported by the Institute for Information & Communications Technology Planning & Evaluation (IITP) grant (No.RS-2022-II221040), and the National Research Foundation of Korea (NRF) grant (No.
@@
 RS-2024-00413957), all of which are funded by the Korean government (MSIT).
+Data availability—The simulation codes and data that support the findings of this article are openly available at [88]
+Appendix
+Figure 5: Energy spectrum $E^{\pm}_{n}$ of the Hamiltonian in Eq.
+(1) as a function of $|v_{r}|-|v_{cr}|$, with $w=5$.
+The spectrum is obtained via numerical diagonalization with a Fock-space cutoff $N_{\max}=3000$, for $\phi=0$ and $\phi=\pi/2$.
+The singular region $0.01\leq|v_{r}|-|v_{cr}|\leq 0.01$, which is excluded from the calculation, is masked out.
+Each line corresponds to an energy level $E_{n}$, and the color indicates its sign.
+Cross markers represent theoretical predictions from Eq.
+(2) for $1\leq n\leq 10$, showing good agreement with the numerical results, except near the singular region.
+When $\phi=\pi/2$, the spectrum displays a distinct feature near the point $|v_{r}|=|v_{cr}|$, as detailed in the main text.
+Figure 6: Energy spectrum $E^{\pm}(x,\phi)$ given in Eq. (4) for $w=1$, $|v|=2$. A periodic array of conical Dirac points appears at $(q,\phi)=(0,m\pi)$ for $m\in\mathbb{Z}$.
+Appendix A: Energy Spectrum—Figure 5 shows the energy spectra obtained via numerical diagonalization of the Hamiltonian in Eq.
+(1) for $\phi=0$ and $\phi=\pi/2$ under $|v_{r}|\neq|v_{cr}|$.
+The analytical expression in Eq.
+(2), indicated by cross markers, shows excellent agreement with the numerical results.
+Notably, the spectra exhibits qualitatively different behaviors near the critical point $|v_{r}|=|v_{cr}|$, depending on the values of $\phi$, as discussed in the main text.
+When $|v_{r}|=|v_{cr}|$, a conical energy spectrum emerges in a (1+1)(1\!+\!1)-dimensional space spanned by $x$ and $\phi$, as described by Eq.
+(4).
+Figure 6 illustrates the resulting periodic Dirac-cone structure in the $(x,\phi)$ space.
+Figure 7: Phase-space trajectories with fixed $\phi_{r}$ and $\phi_{cr}$ swept over $0,2\pi]$.
+Other conditions are the same to Fig.
+2.
+Red solid (blue dashed) curves correspond to the regime $|v_{r}|>|v_{cr}|$ ($|v_{r}|<|v_{cr}|$).
+The geometric properties remain the same, except the condition for whether the trajectory encircles the origin is reversed.
+Thus, the two topological phases can still be distinguished by the phase-space winding number.
+Appendix B: Phase-Space Trajectory—All eigenstates in Eq. (2) are centered at
+$\alpha_{r}(\phi_{r},\phi_{cr})=\frac{w(v_{cr}^{*}-v_{r})}{|v_{r}|-|v_{cr}|^{2}},$ (7)
+where $\Re\{\alpha_{r}\}=\sqrt{2}\langle\hat{x}\rangle$ and $\Im\{\alpha_{r}\}=\sqrt{2}\langle\hat{p}\rangle$.
+Here, $\hat{x}$ and $\hat{p}$ are the position and momentum operators in the ladder basis of $\hat{a}$ and $\hat{a}^{\dagger}$, respectively.
+The additional degree of freedom $\phi_{cr}$ can be eliminated by moving to a rotated frame, $\alpha(\phi)=\alpha_{r}(\phi)e^{i\phi_{cr}}$.
+In this frame, the phase-space center becomes
+$\alpha(\phi)=\frac{w(|v_{cr}|-|v_{r}|e^{i2\phi})}{|v_{r}|^{2}-|v_{cr}|^{2}}.$ (8)
+Fixing $\phi_{cr}$ and varying $\phi_{r}$ from 0 to $2\pi$, the trajectories in Fig.
+2 correspond to the positions of $\alpha(\phi)$ over $\phi\in[0,\pi]$ in the complex $\alpha$-plane.
+Since $\phi_{cr}$ is fixed, we have $\phi\propto-\theta$ for $2\theta=\phi_{cr}-\phi_{r}$, allowing the trajectories to be interpreted as geometric features of the eigenstates as the phase difference between the hopping terms $v_{r}$ and $v_{cr}$ varies from 0 to $2\pi$.
+A single loop of $\alpha(\phi)$ over $\phi\in[0,\pi]$ thus captures the helical structure of the eigenstates along the $\theta$-axis induced by $\hat{R}(\theta)$.
+Alternatively, one may fix $\phi_{r}$ instead of $\phi_{cr}$.
+However, this choice does not affect the global property—whether the phase-space trajectories encircle the origin depends only on the relative values of $|v_{r}|$ and $|v_{cr}|$, as shown in Fig.
+7.
+That is, the choice of fixed parameter determines only the sign of the relation between $\phi$ and $\theta$, while the geometric structure remains unchanged.
+Experimentally, the phase-space trajectories can be measured by preparing an eigenstate and tracking its phase-space distribution as a function of $\phi_{r}$ or $\phi_{cr}$, while keeping the other fixed.
+The choice of which phase parameter to vary depends solely on the experimentalist’s preference.
+The choice of phase parameter to vary is largely a matter of convenience and has no significant impact on the result.
+The phase-space winding number can be calculated as
+$W$ $=\frac{1}{2\pi i}\int_{0}^{\pi}d\phi\,\frac{d}{d\phi}\ln\alpha(\phi)=\frac{1}{2\pi i}\oint_{\mathcal{C}^{\prime}}\frac{1}{t}\,dt,$
+where $t=|v_{cr}|-|v_{r}|e^{2i\phi}$.
+Here, $\mathcal{C}^{\prime}$ is the trajectory traced by $t$ as $\phi$ varies from 0 to $2\pi$.
+According to the residue theorem, the last integral evaluates to $2\pi i$ if the contour $\mathcal{C}^{\prime}$ encloses the origin, and to zero otherwise.
+Therefore, the winding number is given by
+$2W=1-\mathrm{sgn}(|v_{cr}|-|v_{r}|).$
+If instead we fix $\phi_{r}$ and define $\alpha(\phi)=\alpha_{r}(\phi)e^{-i\phi_{r}}$, the contour $\mathcal{C}^{\prime}$ becomes the trajectory traced by $t(\phi)=|v_{cr}|e^{-2i\phi}-|v_{r}|$, resulting in a winding number:
+$2W=1+\mathrm{sgn}(|v_{cr}|-|v_{r}|).$
+Figure 8: Fock-state populations of eigenstates at $w=v_{r}=1$ and $v_{cr}=0.25$.
+For clarity, the weaker hopping $v_{cr}$ is not shown.
+Solid and dashed lines indicate the stronger and weaker of the intra-cell $w$ and inter-cell $v_{r}$ couplings.
+Each population $|\langle n,A|\psi\rangle|^{2}$ for $A\in\{e,g\}$ is encoded by the opacity of the circles, with higher opacity indicating larger population.
+A domain wall appears at $n=1$, and the defect $\left|1,g\right\rangle$ is highlighted by a green box.
+The zero-energy defect state $\left|\psi_{0}\right\rangle$ is localized near the defect and resides on the $\left|g\right\rangle$ sublattice.
+In contrast, the populations of other eigenstates $\left|\psi_{k}^{+}\right\rangle$ are more spread over the FSL and occupy both sublattices.
+Appendix C: Fock-state Populations— This section presents the Fock-state populations of the eigenstates $\left|\psi_{0}\right\rangle$ and $\left|\psi_{k}^{+}\right\rangle$ ($k\in\{1,2,3\}$) with parameters $v_{r}=w=1$ and $v_{cr}=0.25$ used in the numerical simulations of Fig.
+3.
+In this case, a domain wall appears at site $n=1$, with the defect located at $\left|1,g\right\rangle$.
+Figure 8(a) shows the Fock-state population of the zero-energy defect state $\left|\psi_{0}\right\rangle$, which is localized near the defect, indicated by a green box.
+The defect state resides on the sublattice $\left|g\right\rangle$, and the topological protection of its spin polarization $\langle\psi_{0}|\hat{\sigma}_{z}|\psi_{0}\rangle$ prevents excitation into the other sublattice $\left|e\right\rangle$ under parameter noise.
+In contrast, the other eigenstates $\left|\psi_{k}^{+}\right\rangle$ exhibit Fock-state populations extended across the FSL, as shown in Fig.
+8(b)–(d).
+These states occupy both sublattices and are therefore vulnerable to parameter noise, since their spin states are not protected.
+Appendix D: Experimental Realization—The Hamiltonian in Eq.
+(1) can be implemented using trapped ions or superconducting circuits, where the coupling strengths $|v_{r}|$, $|v_{cr}|$, and $w$ correspond to sideband and carrier Rabi frequencies [19, 44, 45, 46, 47, 48].
+Since the energy gap between $E_{0}$ and $E_{\pm 1}$ remains finite unless $|v_{r}|=|v_{cr}|$, the zero-energy defect state can be adiabatically prepared by first activating one of the sidebands ($v_{r}$ or $v_{cr}$) and then gradually increasing both $w$ and the remaining sideband strength.
+Geometric features, such as phase-space trajectories and distributions, can be estimated trough quantum state tomography [71, 69, 70].
+The main limitation of this adiabatic protocol lies in the trade-off between the adiabatic timescale and the coherence time of the zero-energy state.
+The adiabatic timescale is determined by the energy gap $|E_{\pm 1}-E_{0}|=\sqrt{||v_{r}|-|v_{cr}||}$, and thus the coherence time must exceed this timescale to ensure high-fidelity state preparation and measurement.
+In trapped-ion systems, typical red- and blue-sideband Rabi frequencies are on the order of tens of kHz, and $v_{r}$, $v_{cr}$, and $w$ can be independently tuned via radio-frequency (RF) amplitude modulation.
+The motional coherence time near the ground state can approach one second [95], making trapped-ion platforms promising candidates for probing phase-space characteristics through an adiabatic process.
+However, the coherence time is significantly shorter for highly excited Fock states, and the accessible parameter regime may be constrained by the Lamb–Dicke condition [44], which limits both the maximum achievable sideband coupling strength and the Fock-state population.
+In superconducting circuits, sideband Rabi frequencies can reach tens of MHz [48], allowing for faster adiabatic simulations.
+However, the motional coherence time is relatively shorter than in trapped-ion systems, reaching up to tens of milliseconds [96, 97].
+The effective parameter regimes remain similar to trapped-ion systems, as the system’s topological features are determined by relative coupling strengths rather than absolute values.
+In both platforms, the phase factors $\phi_{r}$ and $\phi_{cr}$ are tunable by controlling the relative phases of RF electronics used for sideband generation, enabling precise control of the desired phase-space geometry.
+These systems offer precise control over the drive parameters, making them suitable for the experimental demonstration of single-atom topological features via adiabatic preparation and manipulation of defect states.